# Spatial Adaptation Layer: Interpretable Domain Adaptation For Biosignal Sensor Array Applications

### v1

@@ III Experimental Analysis
LogReg 0.962 0.973 0.972 0.969
CpgmNet 0.884 0.930 0.905 0.935
TABLE II: Intersession Experiments (Test Accuracy)
BN BN + Drpt. LBN LBN+Drpt.
CSL
LogReg 0.147(0.126) 0.227(0.190) - -
LogReg(fine-tune) 0.427(0.326) 0.414(0.266) - -
LogReg(SAL) 0.337(0.274) 0.571(0.339) 0.555(0.404) 0.698(0.394)
CpgmNet 0.151(0.127) 0.184(0.154) - -
CpgmNet(fine-tune) 0.700(0.502) 0.686(0.405) - -
CpgmNet(SAL) 0.386(0.305) 0.604(0.425) 0.133(0.114) 0.721(0.472)
Capgmyo
LogReg 0.439(0.436) 0.447(0.444) - -
LogReg(fine-tune) 0.761(0.752) 0.787(0.728) - -
LogReg(SAL) 0.614(0.606) 0.656(0.550) 0.723(0.714) 0.737(0.638)
CpgmNet 0.401(0.318) 0.402(0.328) - -
CpgmNet(fine-tune) 0.813(0.586) 0.745(0.476) - -
CpgmNet(SAL) 0.582(0.414) 0.535(0.375) 0.652(0.430) 0.587(0.391)
TABLE III: Ablation Experiments (Test Accuracy)
LBN          (no affine) No translation Translation only Circumferential translation only
CSL
LogReg 0.433(0.253) 0.497(0.281) 0.647(0.369) 0.644(0.363)
CpgmNet 0.526(0.308) 0.592(0.357) 0.723(0.468) 0.714(0.454)
Capgmyo
LogReg 0.553(0.495) 0.628(0.544) 0.709(0.621) 0.704(0.62)
CapgmNet 0.492(0.339) 0.530(0.351) 0.577(0.367) 0.607(0.387)
III-A Experiment 1: Intrasession Performance
The first experiment involves obtaining the intrasession performance scores for the aforementioned classifiers on each of the datasets, acting as a practical upper-bound for intersession performance. As shown in Table I accuracies are high across models, with the optimal condition for each model being above 93% accuracy. Interestingly, even a logistic regressor is able to achieve competitive performances, outperforming CapgmyoNet with Capgmyo data.
III-B Experiment 2: Intersession Performance
Intersession performance is measured as the accuracy from training on one session and testing on another session of the same subject. In our approach, SAL and LBN are optimized during an adaptation step, which is compared against fine-tuning. We expected that since the data from Capgmyo is not a regular grid (variable interelectrode distance), assumptions made by SAL would distort the images and not be as beneficial towards performance, whereas the regular grids used in CSL make it the ideal SAL test case. Performances reported use majority voting over the full active segment for CSL and 150ms window for Capgmyo, as in [5]. Non majority-voted performances are reported in parentheses.
From Table II, as expected, while all models achieved higher performance with fine-tuning for Capgmyo, for the regular grid data in CSL the optimal performance was achieved with SAL, despite using orders of magnitude less parameters. Furthermore, when using SAL, it can be clearly seen across models that the use of LBN significantly improves performance over traditional BN, highlighting the flexibility of learning the optimal baseline activity to subtract over estimating it directly from the data. Surprisingly, input dropout proved vital when using SAL, consistently resulting in accuracy improvements of over 20%, up to an astonishing improvement of almost 60% improvement for CapgmyoNet. It is likely that input dropout significantly probes the model to depend on different input combinations, reducing its likelihood of overfitting to specific channels. Additionally, while SAL leaves the original learned representation intact, further experiments showed that after fine-tuning, performance on the original training set was less than 1% higher than zero-shot intersession, indicative of catastrophic forgetting [10]. Finally, while the CNNs outperform logistic regression, it is promising to see that with SAL applying domain adaptation at the input level, even traditional, interpretable classifiers can achieve comparable performance to SOTA EMG-based classifiers.
III-C Experiment 3: Model Ablation & Interpretability Assessment
Given the results from Experiment 2, we explore the importance of different affine coefficients on the final performance for each model and dataset, using input dropout and LBN. Given that the signal array geometry and often orientation are consistent across experiments in both datasets considered, especially in sEMG, we expected that the translation parameters would be the most significant out of the 7 learnable coefficients. In addition, since electrode displacements along the circumference of the arm are most detrimental to sEMG classification performance as muscles are mostly distributed along the length of the forearm [6, 9], we expected that corrections of such translations with SAL are likely to yield the largest performance improvements.
Thus, we run an ablation experiment of SAL with four different conditions: LBN (no affine), frozen translation coefficients, all SAL coefficients frozen except for translation, and keeping only the circumferential shift parameter learnable. As in Experiment II, performances reported use majority voting over the full active segment for CSL and 150ms window for Capgmyo, as in [5]. Non majority-voted performances are reported in parentheses.
As shown in Table III, while LBN alone accounts for significant improvements, SAL accounts for the majority of the improvements with only 7 learnable parameters. Furthermore, as expected, only keeping 2 learnable translation parameters results in significantly higher performance than the remaining 5 affine parameters. In line with past studies, the performance for only circumferential translation (1 affine coefficient) is nearly identical to the performance using both translation coefficients. Thus, SAL is able to drastically improve performance using few parameters closely associated with the physical cause of performance detriments. While these trends were observed for both datasets, non-translation affine parameters seemed to account for a larger proportion of performance with Capgmyo than CSL, as they are more useful for distortions introduced by the irregular geometry of Capgmyo grids.
IV Conclusion
@@
We have shown that SAL can be prepended to any differentiable model, from a logistic regressor to CNNs, and achieve higher intersession performances than even standard fine-tuning with orders of magnitude less parameters for biosignal arrays. In addition, LBN better accounts for baseline activity fluctuations over traditional BN making it particularly well suited for biosignal sensor arrays such as sEMG. Ablation experiments suggest that learnt SAL coefficients are directly associated with the physical causes of the performance detriment. Interestingly, major improvements in sEMG systems by only accounting for circumferential translations suggest the possibility of applying the proposed approach to more practical 1D sensor arrays, such as the one in [1].
@@
Using one repetition per class, as in this study, is commonplace in sEMG literature [11, 5], with some studies using half [12] or even all [5] of the target data for adaptation. However, this becomes impractical for large number of classes. Thus, given that SAL and LBN apply domain adaptation at an input level, and electrode displacement and baseline fluctuations being invariant to class labels, future studies will explore whether our approach can adapt to one repetition of one (or a few) class(es), unlike present domain adaptation approaches.

### v2

@@ IV-D Model Ablation & Interpretability Assessment
LogReg 0.553 0.628 0.709 0.704
CapgmNet 0.492 0.530 0.577 0.607
V Conclusion
@@
We have shown that SAL can be prepended to any differentiable model, from a logistic regressor to CNNs, and achieved higher intersession performances than even standard fine-tuning with orders of magnitude less parameters for biosignal arrays. In addition, LBN better accounted for baseline activity fluctuations over traditional BN making it particularly well suited for biosignal sensor arrays such as sEMG. Ablation experiments suggest that learned SAL coefficients are directly associated with the physical causes of the performance detriment. Interestingly, major improvements in sEMG systems by only accounting for circumferential translations suggest the possibility of applying the proposed approach to more practical 1D sensor arrays, such as the one in [1].
@@
Using one repetition per class, as in this study, is commonplace in sEMG literature [11], with some studies using half [12] or even all [5] of the target data for adaptation. However, this becomes impractical for large number of classes. Thus, given that SAL and LBN apply domain adaptation at an input level, and electrode displacement and baseline fluctuations being invariant to class labels, future studies will explore whether our approach can adapt to one repetition of one (or a few) class(es), unlike present domain adaptation approaches.